Using Maxwell's equations in free space,derive the wave equationn for the z- component of the electric field vector.
1
Expert's answer
2019-04-01T10:34:04-0400

The Maxwell's equations in free space can be written as follows:


"\\nabla \\times \\vec{E} = - \\frac{1}{c} \\frac{\\partial \\vec{H}}{\\partial t}, \\quad (1) \\\\\n\\nabla \\times \\vec{H} = \\frac{1}{c} \\frac{\\partial \\vec{E}}{\\partial t}\t+ \\frac{4 \\pi}{c} \\vec{j}, \\quad (2)\\\\\n\\nabla \\cdot \\vec{E} = \\rho, \\quad (3) \\quad\\quad \\nabla \\cdot \\vec{H} = 0. \\quad (4)"

Taking the curl operation from the first equation, we obtain:


"\nabla \times (\nabla \times \vec{E}) = \nabla \times \left( - \frac{1}{c} \frac{\partial \vec{H}}{\partial t} \right)"

This expression can be simplified as follows:


"\nabla(\nabla \cdot \vec{E}) - \Delta \vec{E} = - \frac{1}{c} \frac{\partial }{\partial t} \left( \nabla \times \vec{H} \right)"

After substitution of the corresponding expressions from the 2nd and 3rd Maxwell's equations, we derive:


"\nabla \rho - \Delta \vec{E} = - \frac{1}{c^2} \frac{\partial^2 \vec{E}}{\partial t^2} - \frac{4 \pi}{c^2} \frac{\partial \vec{j}}{\partial t}"

This expression can be re-written in the following way:


"\Delta \vec{E} - \frac{1}{c^2} \frac{\partial^2 \vec{E}}{\partial t^2} = \nabla \rho + \frac{4 \pi}{c^2} \frac{\partial \vec{j}}{\partial t}"

In order to derive the expression for the z-component of the electric field vector, one should project this expression onto the z-axis:


"\Delta E_z - \frac{1}{c^2} \frac{\partial^2 E_z}{\partial t^2} = \frac{\partial \rho}{\partial z} + \frac{4 \pi}{c^2} \frac{\partial j_z}{\partial t},"

which is the answer on the question.